### v1

@@ 3.2 Baselines models
To evaluate the performance of our model, we considered two groups of baseline models. The first group consists of supervised learning methods, including MoleculeNet36, D-MPNN37, ML-MPNN38, FP-GNN23, MVGNN39, LineEvo40, HimGNN41. The second group is based on self-supervised learning approaches, including GROVER6, MGSSL25, HiMOL42. A brief introduction to the aforementioned baseline model is provided below.
  • MoleculeNet36 employs a diverse set of benchmark models for molecular property prediction, encompassing traditional machine learning algorithms such as Random Forest, as well as Graph-based approaches including GCN.
  • D-MPNN37 utilizes a directed edge-based message passing framework to more effectively capture information about chemical bonds, which enhances the performance of molecular property prediction.
  • ML-MPNN38 enhances molecular property prediction by integrating sequence-level information in addition to conventional graph-level features
  • FP-GNN23 aggregates molecular fingerprint features and graph-based representations to facilitate molecular property prediction.
  • MVGNN39 introduces a novel Cross-Dependent Message Passing Scheme that facilitates mutual interaction between node and edge features, thereby improving the model’s expressive power.
  • LineEvo40 proposes a new architectural component, the LineEvo Layer, which captures hierarchical granular features from molecular structures to enhance prediction performance.
  • HimGNN41 introduces motif-level representations to extract structural information from molecular graphs, which are subsequently integrated with traditional graph features to improve the accuracy of molecular property prediction.
@@
  • MGSSL25 enhances molecular property prediction by combining motif-level generative tasks with multi-level self-supervised pretraining on molecular graphs
  • HiMOL42 integrates node-level, motif-level, and graph-level representations, leveraging a pretraining framework to enhance performance on molecular predictions tasks.
The performance metrics of the baseline models were obtained from existing literature. Specifically, for most datasets, we adopted the baseline results reported in the HimGNN41, while for datasets under the PDBbind category, the results were referenced from the FP-GNN23. To ensure a fair comparison with these baseline models, we aligned certain experimental settings with those used in the corresponding studies.
@@ 3.3 Implementation details
We implemented our model using the RDKit and Pytorch, employing the Adam 43 optimizer to update model parameters. Hyperparameter tuning was conducted using Bayesian Optimization 44 to efficiently explore the search space. Specifically, we optimized seven key hyperparameters: (1) the number of Transformer layers $N$, (2) the number of attention heads $h$, (3) the dimensionality of each attention head $d_{k}$, (4) the output dimension of the GAT module $G_{out}$, (5) the dropout rate in the feedforward network $F_{d}$, (6) the dropout rate in the GAT module $G_{d}$, and (7) the dropout rate in the Transformer module $A_{d}$. The results of hyperparameter search are summarized in Table 3. For each hyperparameter configuration, the model was trained using 5 different random seeds, and the configuration with the highest average performance was selected as the final setting. Dataset splitting was performed according to the protocol described in Table 2. In the case of random splits, the dataset was divided into training, validation, and test sets with a ratio of 8:1:1. The validation set was used to select the final model for the better generalization ability. For each dataset, we trained our model using 10 different random seeds. All models were trained on NVIDIA GeForce RTX 4090.
Iter $N$ $h$ $d_{k}$ $G_{out}$ $G_{d}$ $F_{d}$ $A_{d}$ RMSE
@@ 3.4 Performance analysis
To evaluate the effectiveness of our proposed model, we conduct a comprehensive comparison several strong baselines across a suite of benchmark datasets covering both classification and regression tasks. Table 4 presents the mean and standard deviation of ROC-AUC scores on the classification benchmarks. The proposed model achieves the highest performance on three out of five datasets and ranks second on the remaining two, demonstrating its robustness and consistency across diverse molecular property prediction tasks. Similarly, Table 5 and Table 6 summarize the RMSE results for regression tasks. Our model consistently outperforms all baselines, highlighting its superior ability to capture intrinsic molecular features and the complex relationships among atoms and substructures. In summary, the experimental results on both classification and regression tasks demonstrate the strong generalization ability and effectiveness of our model. By leveraging multi-level structural information and a hierarchical representation learning framework, the proposed model is capable of capturing complex molecular features and dependencies, leading to consistently superior performance across a variety of molecular property prediction benchmarks. These findings validate the design of our approach and its potential for broad applicability in computational chemistry and drug discovery tasks.
@@
Table 4: Performance on classification tasks in terms of ROC-AUC. Higher is better and the best results in bold.
@@
Table 5: Performance on PDBbind datasets in terms of RMSE. Lower is better and the best results in bold.
@@
Table 6: Performance on regression tasks in terms of RMSE. Lower is better. Best results in bold.
@@ 3.5 Ablation study
In order to access the contribution of each component within our model architecture, we performed a series of ablation experiments targeting the Graph Transformer, the Molecular Graph representation, and the Fingerprint-based features. The ablation study results are presented in Figure 3, Figure 4, Figure 5.
@@
Figure 4 illustrates the results of the ablation experiments focusing on the Fingerprint module. We replaced the original concatenation of three molecular fingerprints in our model with two alternative configurations: using only the Morgan fingerprint and using the three fingerprints adopted in FP-GNN23(MACCS, PubChem, and ErG). Additionally, we performed an ablation in which the entire Fingerprint component was removed. The consistent superiority of the Morgan, PubChem, ErG fingerprints combination across all datasets demonstrates the effectiveness of combining diverse and complementary fingerprint features. In particular, the inclusion of Morgan fingerprints provides substantial gains over the MACCS-based combination used in prior models such as FP-GNN, suggesting that Morgan captures more informative substructural details. Moreover, the performance degradation observed when removing the Fingerprint component altogether underscores its importance in supplementing graph-based representation, enabling the model to leverage complementary molecular information for more accurate predictions.
@@ 3.6 Visualization analysis
To gain a deeper understanding of the contributions of different components in MLFGNN, we conducted interpretability analysis using SHAP (SHapley Additive exPlanation) values. SHAP is a game-theoretic approach that explains the output of machine learning models by assigning each feature an importance value for a particular prediction. It is based on Shapley values from cooperative game theory and ensure consistency and local accuracy in feature attribution 45.
Specifically, we calculated SHAP values separately for the fingerprint component and the molecular graph component, enabling us to quantify and compare their respective contributions to the model’s predictions. As shown in Figure 6, the contribution analysis across various datasets reveals that the fingerprint and molecular graph components exhibit varying levels of importance depending on the specific prediction task. In some cases, the molecular graph component plays a more dominant role, while in others, the fingerprint component contributes more significantly. This variation indicates that the two components capture different but complementary aspects of molecular information. Overall, the results suggest that combining both representations enables the model to better generalize across diverse tasks by leveraging the strengths of each modality.
To further understand the respective contributions of the Graph Attention Layer and the Graph Transformer Layer, which correspond to local and global information respectively, we analyzed the learned weighting coefficients from the Adaptive Mixture Layer. These coefficients indicate the relative influence of each component in forming the final molecular representation.
As illustrated in Figure 7, the contributions ratios between the Graph Attention and Graph Transformer components vary across datasets. This variation indicates that local and global information provide complementary insights, enabling the model to adaptively extract features that are well-suited to requirements of downstream tasks. These finding further support the effectiveness of incorporating both components in molecular representation learning.
@@
Figure 8 and Figure 9 illustrate the molecular structures alongside atom-level feature correlation heatmaps at different stages of MLFGNN. The visualizations highlight how atomic relationships evolve before training, after the GAT layer, and after the Mixed-information layer. The heatmaps reveal how different components of the model contribute to refinement of atomic feature representations. After passing through the GAT layer, localized structural patterns such as rings and specific functional groups, are more prominently captured, reflecting the GAT’s strength in modeling local dependencies. When the Graph Transformer is further integrated integrated via the Adaptive Mixture Layer, the resulting representations incorporate additional long-range dependencies. This enables the model to better address challenges like activity cliffs, where subtle, spatially distant atomic differences significantly affect molecular properties. Overall, these visualizations demonstrate the complementary roles of local and global information in generating expressive molecular representations.
@@
Figure 10 and Figure 11 provide visualizations of atom-level contributions to molecular graph representations for two prediction tasks. In both cases, the model assigns higher importance to chemically meaningful substructures. For instance, in Figure 10, the highlighted atoms correspond to function groups, such as aromatic rings and nitrogen-containing motifs, which are known to influence BBB permeability. Similarly, in Figure 11, atoms within hydrophobic aromatic rings of polar functional groups receive higher attention, aligning well with known determinants of lipophilicity. These observations suggest that the model is capable of capturing domain-relevant chemcial knowledge, supporting the interpretability and reliability of its learned representations.
@@
(a) Visualization of atomic correlation patterns in the BBBP dataset
(b) Visualization of atomic correlation patterns in the ClinTox dataset
(c) Visualization of atomic correlation patterns in the Tox21 dataset
Figure 8: Molecular structures and atom feature heatmaps before and after GAT/mixed layers in some classification datasets. From left to right, each group of images shows: (1) the molecular structure, (2) the atom feature heatmap before training, (3) the atom feature heatmap after the GAT layer, and (4) the atom feature heatmap after the Mixed-information layer
(a) Visualization of atomic correlation patterns in the ESOL dataset
(b) Visualization of atomic correlation patterns in the PDB-R dataset
(c) Visualization of atomic correlation patterns in the Lipo dataset
Figure 9: Molecular structures and atom feature heatmaps before and after GAT/mixed layers in some regression datasets. From left to right, each group of images shows: (1) the molecular structure, (2) the atom feature heatmap before training, (3) the atom feature heatmap after the GAT layer, and (4) the atom feature heatmap after the Mixed-information layer.
@@ 4 CONCLUSIONS
In this paper, we present MLFGNN, a Multi-Level Fusion Graph Neural Network designed to enhance molecular property prediction by combining structural and multi-modal information. Our model addresses the limitations of conventional GNNs by integrating both local and global molecular features via a hybrid GAT–Graph Transformer architecture. Additionally, molecular fingerprints are incorporated through a cross-attention mechanism, enabling adaptive feature selection and improved alignment with downstream tasks. Experimental results across various benchmark datasets validate the effectiveness and generalizability of our approach. Moreover, interpretability analysis confirms that MLFGNN can identify chemically meaningful patterns, including functional groups and long-range dependencies, that are crucial for accurate prediction. Overall, our framework demonstrates the potential of combining intra-graph and inter-modal fusion strategies for robust and interpretable molecular representation learning.
@@ 6.1 Corresponding Author
Hou-biao Li — School of Mathematical Sciences, University of Electronic Science and Technology of China, Chengdu 610054, China; https://orcid.org/$0$-0002-7268-307X; Email: lihoubiao0189@163.com.

### v2

@@ 3.2 Baselines models
To evaluate the performance of our model, we considered two groups of baseline models. The first group consists of supervised learning methods, including MoleculeNet40, D-MPNN41, ML-MPNN42, FP-GNN30, MVGNN43, LineEvo44, HimGNN45, AttentiveFP37, ResGAT46. The second group is based on self-supervised learning approaches, including GROVER6, MGSSL33, HiMOL47, MolTailor48, MolGraph-LarDo49, PremuNet22, MolGT23. A brief introduction to the aforementioned baseline model is provided below.
  • MoleculeNet40 employs a diverse set of benchmark models for molecular property prediction, encompassing traditional machine learning algorithms such as Random Forest, as well as Graph-based approaches including GCN.
  • D-MPNN41 utilizes a directed edge-based message passing framework to more effectively capture information about chemical bonds, which enhances the performance of molecular property prediction.
  • ML-MPNN42 enhances molecular property prediction by integrating sequence-level information in addition to conventional graph-level features
  • FP-GNN30 aggregates molecular fingerprint features and graph-based representations to facilitate molecular property prediction.
  • MVGNN43 introduces a novel Cross-Dependent Message Passing Scheme that facilitates mutual interaction between node and edge features, thereby improving the model’s expressive power.
  • LineEvo44 proposes a new architectural component, the LineEvo Layer, which captures hierarchical granular features from molecular structures to enhance prediction performance.
  • HimGNN45 introduces motif-level representations to extract structural information from molecular graphs, which are subsequently integrated with traditional graph features to improve the accuracy of molecular property prediction.
@@
  • MGSSL33 enhances molecular property prediction by combining motif-level generative tasks with multi-level self-supervised pretraining on molecular graphs.
  • HiMOL47 integrates node-level, motif-level, and graph-level representations, leveraging a pretraining framework to enhance performance on molecular predictions tasks.
  • AttentiveFP37 incorporates a new graph attention mechanism to learn molecular representations directly from molecular graphs.
  • ResGAT46 introduces a residual graph attention network that integrates multi-head attention with residual connections to enhance message passing in molecular graphs.
  • MolTailor48 proposes a prompt-based framework that generates task-specific molecular representations by conditioning a pre-trained language model on natural language task descriptions.
  • MolGraph-LarDo49 presents a hybrid framework that combines large language models with lightweight, domain-specfic graph neural networks to enhance molecular representation learning.
  • PremuNet22 introduces a pre-trained multi-representation fusion network that integrates graph-based, sequence-based, and descriptor-based molecular representations.
  • MolGT23 proposes a multimodal pretraining framework for molecular representation that integrates graph structure, SMILES sequences, and expert knowledge using a unified graph transformer architecture.
The performance metrics of the baseline models were obtained from existing literature. Specifically, for most datasets, we adopted the baseline results reported in HimGNN 45, while for the PDBbind datasets, results were referenced from FP-GNN 30. For other baseline models, we used the results as reported in their original papers, including ResGAT46, MolTailor48, MolGraph-LarDo49, PremuNet22, and MolGT23. To ensure a fair comparison with these baseline models, we aligned certain experimental settings with those used in the corresponding studies.
@@ 3.3 Implementation details
We implemented our model using the RDKit and Pytorch, employing the Adam 50 optimizer to update model parameters. Hyperparameter tuning was conducted using Bayesian Optimization 51 to efficiently explore the search space. Specifically, we optimized seven key hyperparameters: (1) the number of Transformer layers $N$, (2) the number of attention heads $h$, (3) the dimensionality of each attention head $d_{k}$, (4) the output dimension of the GAT module $G_{out}$, (5) the dropout rate in the feedforward network $F_{d}$, (6) the dropout rate in the GAT module $G_{d}$, and (7) the dropout rate in the Transformer module $A_{d}$. The results of hyperparameter search are summarized in Table 3. For each hyperparameter configuration, the model was trained using 5 different random seeds, and the configuration with the highest average performance was selected as the final setting. To enhance the generalizability of the model across datasets, we conducted hyperparameter optimization on the Lipophilicity dataset, which features a chemically diverse set of molecules with a broad range of sizes and structural complexities. As a regression task with continuous-valued targets and rich molecular variation, Lipophilicity serves as a strong candidate for identifying robust hyperparameter configurations. The optimal settings derived from this dataset were then directly applied to other benchmark datasets without additional tuning. As shown in Table 4, Table 5 and Table 6, the transferred configurations yielded consistently competitive results such as the Transformer depth, GAT output dimension, and dropout rates exhibit strong transferability when optimized on a representative and heterogeneous dataset. These findings indicate that Lipophilicity-based tuning can provide a reliable initialization for related molecular property prediction tasks, thereby reducing the computational overhead of dataset-specific hyperparameters searches. For data splitting, we followed the protocol adopted in Wu et al.40. Specifically, MoleculeNet utilizes two primary splitting strategies: random splitting, which randomly partitions molecules into training, validation, and test sets, and scaffold splitting, which separates molecules based on their Bemis-Murcko scaffolds52 to assess model generalization to novel chemical structures. Following this convention, we split each dataset according to the MoleculeNet-recommended strategy using an 8:1:1 ratio for training, validation, and testing. The validation set was used to select the final model for the better generalization ability. For each dataset, we trained our model using 10 different random seeds. All models were trained on NVIDIA GeForce RTX 4090.
Iter NN $h$ $d_{k}$ $G_{out}$ $G_{d}$ $F_{d}$ $A_{d}$ RMSE
@@ 3.4 Performance analysis
To evaluate the effectiveness of our proposed model, we conduct a comprehensive comparison several strong baselines across a suite of benchmark datasets covering both classification and regression tasks. Table 4 presents the mean and standard deviation of ROC-AUC scores on the classification benchmarks. The proposed model achieves the highest performance on two out of five datasets and ranks second on one, demonstrating its robustness and consistency across diverse molecular property prediction tasks. Similarly, Table 5 and Table 6 summarize the RMSE results for regression tasks. Our model consistently outperforms all baselines, highlighting its superior ability to capture intrinsic molecular features and the complex relationships among atoms and substructures. In summary, the experimental results on both classification and regression tasks demonstrate the strong generalization ability and effectiveness of our model. By leveraging multi-level structural information and a hierarchical representation learning framework, the proposed model is capable of capturing complex molecular features and dependencies, leading to consistently superior performance across a variety of molecular property prediction benchmarks. These findings validate the design of our approach and its potential for broad applicability in computational chemistry and drug discovery tasks.
@@
AttentiveFP 0.810 ± 0.050 0.881 ± 0.040 0.640 ± 0.060 0.860 ± 0.050 0.845 ± 0.020
ResGAT 0.822 ± 0.030 0.875 ± 0.040 0.630 ± 0.050 0.839 ± 0.050 0.888 ± 0.010
MolTailor 0.811 ± 0.021 0.806 ± 0.013 0.923 ± 0.044
MolGraph-LarDo 0.830 ± 0.011 0.731 ± 0.015 0.609 ± 0.023
PremuNet 0.843 ± 0.005 0.733 ± 0.006 0.626 ± 0.008 0.740 ± 0.012 0.992 ± 0.002
MolGT 0.845 ± 0.001 0.737 ± 0.010 0.654 ± 0.004 0.758 ± 0.002 0.889 ± 0.004
@@
Table 4: Performance on classification tasks in terms of ROC-AUC. Higher values indicate better performance. The best results are highlighted in bold, and the second-best results are underlined. "–" means data unavailable.
@@
Table 5: Performance on PDBbind datasets in terms of RMSE. Lower values indicate better performance. The best results are highlighted in bold
@@
AttentiveFP 1.522 ± 0.130 3.558 ± 0.720 1.123 ± 0.030
ResGAT 0.812 ± 0.070 1.473 ± 0.310 0.683 ± 0.020
MolTailor 0.723 ± 0.049 1.788 ± 0.326 0.810 ± 0.016
MolGraph-LarDo 1.276 ± 0.053 2.356 ± 0.078 0.748 ± 0.017
PremuNet 0.851 ± 0.038 1.858 ± 0.205 0.730 ± 0.010
MolGT 0.839 ± 0.006 0.788 ± 0.013
@@
Table 6: Performance on regression tasks in terms of RMSE. Lower values indicate better performance. The best results are highlighted in bold, and the second-best results are underlined. "–" means data unavailable.
@@ 3.5 Ablation study
In order to access the contribution of each component within our model architecture, we performed a series of ablation experiments targeting the Graph Transformer, the Molecular Graph representation, and the Fingerprint-based features. The ablation study results are presented in Figure 2, Figure 3, Figure 4. Additionally, we conducted further ablation experiments to explore more detailed architectural variations. These supplementary results are provided in the Supporting Information for completeness and reference.
@@
Figure 3 illustrates the results of the ablation experiments focusing on the Fingerprint module. We replaced the original concatenation of three molecular fingerprints in our model with two alternative configurations: using three single fingerprints and using the three fingerprints adopted in FP-GNN30(MACCS, PubChem, and ErG). Additionally, we performed an ablation in which the entire Fingerprint component was removed. The consistent superiority of the Morgan, PubChem, ErG fingerprints combination across all datasets demonstrates the effectiveness of combining diverse and complementary fingerprint features. In particular, the inclusion of Morgan fingerprints provides substantial gains over the MACCS-based combination used in prior models such as FP-GNN, suggesting that Morgan captures more informative substructural details. Moreover, the performance degradation observed when removing the Fingerprint component altogether underscores its importance in supplementing graph-based representation, enabling the model to leverage complementary molecular information for more accurate predictions.
@@ 3.6 Visualization analysis
3.6.1 Analysis of Model Architecture Contributions
To gain a deeper understanding of the internal mechanisms of MLFGNN, we first analyzed the contributions of its architectural components through interpretability techniques and internal weight inspections. We employed SHAP (SHapley Additive exPlanation) values53 to evaluate the relative contributions of the fingerprint-based and molecular graph-based representations. SHAP is a game-theoretic approach for explaining model outputs by attributing importance values to individual input features. By computing SHAP values separately for the fingerprint and graph components, we were able to quantify their respective roles in the final prediction. As illustrated in Figure 5 , the contribution levels vary across datasets. Specifically,some tasks rely more heavily on molecular graphs, while others benefit more from fingerprints. This variation confirms that the two components provide complementary information, and their combination enhances the model’s generalization ability across diverse molecular property prediction tasks.
To further dissect the architectural design, we analyzed the Adaptive Mixture Layer, which integrates outputs from the Graph Attention Network (GAT) and the Graph Transformer layers. These two modules are responsible for capturing local and global structural information, respectively. By examining the learned mixing coefficients, we can assess the dynamic weighting between local and global features. As shown in Figure 6, the model adaptively adjusts the relative contributions of GAT and Transformer layers based on dataset characteristics. This confirms the model’s ability to extract task-specific information by balancing localized structural motifs with long-range dependencies, thus validating the necessity of incorporating both modules.
3.6.2 Interpretation of Molecular Structures and Feature Attribution
@@
Figure 7 and Figure 8 illustrate the molecular structures alongside atom-level feature correlation heatmaps at different stages of MLFGNN. The visualizations highlight how atomic relationships evolve before training, after the GAT layer, and after the Mixed-information layer. The heatmaps reveal how different components of the model contribute to refinement of atomic feature representations. After passing through the GAT layer, localized structural patterns such as rings and specific functional groups, are more prominently captured, reflecting the GAT’s strength in modeling local dependencies. When the Graph Transformer is further integrated via the Adaptive Mixture Layer, the resulting representations incorporate additional long-range dependencies. This enables the model to better address challenges like activity cliffs, where subtle, spatially distant atomic differences significantly affect molecular properties. Overall, these visualizations demonstrate the complementary roles of local and global information in generating expressive molecular representations.
@@
Figure 9 and Figure 10 provide visualizations of atom-level contributions to molecular graph representations for two prediction tasks. In both cases, the model assigns higher importance to chemically meaningful substructures. For instance, in Figure 9, the highlighted atoms correspond to function groups, such as aromatic rings and nitrogen-containing motifs, which are known to influence BBB permeability. Similarly, in Figure 10, atoms within hydrophobic aromatic rings of polar functional groups receive higher attention, aligning well with known determinants of lipophilicity. These observations suggest that the model is capable of capturing domain-relevant chemical knowledge, supporting the interpretability and reliability of its learned representations.
@@
(a) Visualization of atomic correlation patterns in the BBBP dataset.
(b) Visualization of atomic correlation patterns in the ClinTox dataset.
(c) Visualization of atomic correlation patterns in the Tox21 dataset.
Figure 7: Molecular structures and atom feature heatmaps before and after GAT/mixed layers in some classification datasets. From left to right, each group of images shows: (1) the molecular structure, (2) the atom feature heatmap before training, (3) the atom feature heatmap after the GAT layer, and (4) the atom feature heatmap after the Mixed-information layer. (a) Visualization of atomic correlation patterns in the BBBP dataset. (b) Visualization of atomic correlation patterns in the ClinTox dataset. (c).Visualization of atomic correlation patterns in the Tox21 dataset.
(a) Visualization of atomic correlation patterns in the ESOL dataset.
(b) Visualization of atomic correlation patterns in the PDB-R dataset.
(c) Visualization of atomic correlation patterns in the Lipo dataset.
Figure 8: Molecular structures and atom feature heatmaps before and after GAT/mixed layers in some regression datasets. From left to right, each group of images shows: (1) the molecular structure, (2) the atom feature heatmap before training, (3) the atom feature heatmap after the GAT layer, and (4) the atom feature heatmap after the Mixed-information layer. (a) Visualization of atomic correlation patterns in the ESOL dataset. (b) Visualization of atomic correlation patterns in the PDB-R dataset. (c) visualization of atomic correlation patterns in the Lipo dataset.
@@ 4 CONCLUSIONS
In this paper, we present MLFGNN, a Multi-Level Fusion Graph Neural Network designed to enhance molecular property prediction by combining structural and multi-modal information. Our model addresses the limitations of conventional GNNs by integrating both local and global molecular features via a hybrid GAT–Graph Transformer architecture. Additionally, molecular fingerprints are incorporated through a cross-attention mechanism, enabling adaptive feature selection and improved alignment with downstream tasks. Experimental results across various benchmark datasets validate the effectiveness and generalizability of our approach. Moreover, interpretability analysis confirms that MLFGNN can identify chemically meaningful patterns, including functional groups and long-range dependencies, which are crucial for accurate prediction. Overall, our framework demonstrates the potential of combining intra-graph and inter-modal fusion strategies for robust and interpretable molecular representation learning.
@@ 6.1 Corresponding Author
Hou-biao Li — School of Mathematical Sciences, University of Electronic Science and Technology of China, Chengdu 610054, China; https://orcid.org/0000-0002-7268-307X; Email: lihoubiao0189@163.com.